# Subgradient Convergence Implies Subdifferential Convergence on Weakly Convex Functions: With Uniform Rates Guarantees

### v1

@@ A.1.1 Chaining Argument
$\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})\leq\max_{b\in H(\xi_{1:m})}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}b_{i}u_{v}(\xi_{i}).$
@@
$V(\tau,\xi_{1:m})\leq\mathbb{E}_{\epsilon_{1:m}}\left[\exp\left(\tau\cdot\max_{b\in H(\xi_{1:m})}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}b_{i}u_{v}(\xi_{i})\right)\right].$
@@
$\begin{split}V(\tau,\xi_{1:m})&\leq\mathbb{E}_{\epsilon_{1:m}}\left[\sum_{b\in H(\xi_{1:m})}\sum_{v\in V_{1/2}}\exp\left(\frac{\tau}{m}\cdot\sum_{i=1}^{m}\epsilon_{i}b_{i}u_{v}(\xi_{i})\right)\right]\\ &=\sum_{b\in H(\xi_{1:m})}\sum_{v\in V_{1/2}}~{}\prod_{i=1}^{m}\mathbb{E}_{\epsilon_{i}}\left[\exp\bigg{(}\frac{\tau}{m}\cdot\epsilon_{i}b_{i}u_{v}(\xi_{i})\bigg{)}\right]\end{split}$ (A.15)
where the symbol $\mathbb{E}_{\epsilon_{i}}$ means the expectation is taken solely over the Rademacher variables $\epsilon_{i}$.
Given that for any $b_{i}\in\{0,1\}$ and $q_{i}\in\mathbb{R}$, the expectation over $\epsilon_{i}$ obeys:
$\mathbb{E}_{\epsilon_{i}}[\exp(\epsilon_{i}b_{i}q_{i})]=\cosh(b_{i}q_{i})\leq\cosh(q_{i})=\mathbb{E}_{\epsilon_{i}}[\exp(\epsilon_{i}q_{i})]$
@@
$\begin{split}V(\tau,\xi_{1:m})&\leq\sum_{v\in V_{1/2}}\sum_{b\in H(\xi_{1:m})}\prod_{i=1}^{m}\mathbb{E}_{\epsilon_{i}}\left[\exp\left(\frac{\tau}{m}\cdot\epsilon_{i}u_{v}(\xi_{i})\right)\right]\\ &=|H(\xi_{1:m})|\sum_{v\in V_{1/2}}\prod_{i=1}^{m}\mathbb{E}_{\epsilon_{i}}\left[\exp\left(\frac{\tau}{m}\cdot\epsilon_{i}u_{v}(\xi_{i})\right)\right].\end{split}$
@@
$V(\tau,\xi_{1:m})\leq(2m)^{{\rm vc}(\mathcal{W})}\sum_{v\in V_{1/2}}\prod_{i=1}^{m}\mathbb{E}_{\epsilon_{i}}\left[\exp\left(\frac{\tau}{m}\cdot\epsilon_{i}u_{v}(\xi_{i})\right)\right].$ (A.16)
@@
$\begin{split}U(\tau)&\leq\mathbb{E}_{\xi_{1:m}}[V(4\tau,\xi_{1:m})]\leq(2m)^{{\rm vc}(\mathcal{W})}\sum_{v\in V_{1/2}}\prod_{i=1}^{m}\mathbb{E}_{\epsilon_{i},\xi_{i}}\left[\exp\left(\frac{4\tau}{m}\cdot\epsilon_{i}u_{v}(\xi_{i})\right)\right]\end{split}$ (A.17)
where $\mathbb{E}_{\epsilon_{i},\xi_{i}}$ denotes the joint expectation taken over $\epsilon_{i}$ and $\xi_{i}$.
Since $u_{v}(\xi)$ is $\bar{\sigma}$-sub-exponential, then $\epsilon_{i}u_{v}(\xi_{i})$ is also $\bar{\sigma}$-sub-exponential for every $v$ by definition. Using [Ver18, Proposition 2.7.1(e)], this implies that, for every $\theta_{2}$, $v$, $\tau\in(0,cm/\bar{\sigma})$:
$\mathbb{E}_{\epsilon_{i},\xi_{i}}\left[\exp\left(\frac{4\tau}{m}\cdot\epsilon_{i}u_{v}(\xi_{i})\right)\right]\leq\exp\left(C\frac{\bar{\sigma}^{2}\tau^{2}}{m^{2}}\right)$
@@
By applying Markov’s inequality under the outer integral [vdVW96] (which follows from the monotone property of outer integral discussed below Definition A.1), we establish that there exists a universal constant $c>0$ such that for every $u\geq 0$:
@@ A.1.3 A Basic Bound
This section proves Lemma A.3. By definition:
@@ A.1.4 A Result on Measurability
We prove Lemma A.5. Let $\mathcal{Z}_{1}^{o}$ and $\mathbb{R}^{o}$ be any countable dense subsets of $\mathcal{Z}_{1}$ and $\mathbb{R}$, respectively. Our main observation is that the supremum over the uncountable sets $\mathcal{Z}_{1},\mathbb{R}$ is equal to the supremum over the countable sets $\mathcal{Z}_{1}^{o},\mathbb{R}^{o}$, holding for every data instances $\xi_{1:m},\epsilon_{1:m}$:
$\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})=\sup_{z_{1}\in\mathcal{Z}_{1}^{o},t\in\mathbb{R}^{o}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})$ (A.19)
Since the RHS is the pointwise supremum over countable measurable functions of $\xi_{1:m},\epsilon_{1:m}$, it is measurable. Therefore, the LHS must also be measurable under the product space. To complete the proof of Lemma A.5, it suffices to prove equation (A.19).
Fix the values of $\xi_{1:m}$ and $\epsilon_{1:m}$. It suffices to show that for every $z_{1}\in\mathcal{Z}_{1}$ and $t\in\mathbb{R}$, there exist $z_{1}^{o}\in\mathcal{Z}_{1}^{o}$ and $t^{o}\in\mathbb{R}^{o}$ such that for every $i=1,2,\ldots,m$:
@@
Recall that $\psi_{z_{1},t}(\xi)=\mathbf{1}\{o(z_{1},\xi)\geq t\}$. First, we note an important observation. For every $z_{1}$, there exists $t^{o}\in\mathbb{R}^{o}$ with $t^{o}<t$ such that for every $i=1,2,\ldots,m$:
@@
This proves that for every $z_{1}\in\mathcal{Z}_{1}$ and $t\in\mathbb{R}$, there is the existence of the pair $z_{1}^{o}\in\mathcal{Z}_{1}^{o}$ and $t^{o}\in\mathbb{R}^{o}$ such that equation (A.20) holds for all $i=1,2,\ldots,m$. This completes the proof of Lemma A.5.
Appendix B Proof of Theorem 5.2
Our approach comprises two main steps. First, we demonstrate that every stationary point of the empirical objective $\Phi_{S}$ must be approximately stationary for the population objective $\Phi$, leveraging the uniform convergence of subdifferentials in Theorem 5.1. Second, we utilize established results from the literature [DDP20, Section 5] on these (approximate) stationary points of $\Phi$ to pinpoint the location of empirical stationary points.
To set up the stage, we recall the location of the stationary point and approximate stationary point of the population objective $\Phi$ formally derived in [DDP20].
Theorem B.1 ([DDP20, Theorem 5.2]).
@@
The following Theorem B.2 can be easily deduced from [DDP20, Theorem 5.2+Theorem 5.3]. Recall that $\mathcal{Z}=\{x:0\in\partial\Phi(x)\}$ denotes the set of stationary points of $\Phi$.
Theorem B.2 ([DDP20, Theorem 5.3]).
@@
$\epsilon={\rm dist}(0;\partial\Phi(x))\leq\gamma\left\|{x}\right\|,$
@@
The following result is immediate given [DDP20, Corollary 6.3].
Proposition B.3 ([DDP20, Corollary 6.3]).
There exist a numerical constant $C>0$ so that for $m\geq Cd$, with an inner probability at least $1-2\exp(-d)$ [^1] The original expression in the paper [DDP20], $1-2\exp(-\min\{c_{2}m,d^{2}\})$, should be corrected to $1-2\exp(-d)$. , the set of stationary points $\mathcal{Z}_{S}$ of $\Phi_{S}$ satisfies:
@@
We shall build on top of these existing results to establish Theorem 5.2. First, Proposition B.3 implies there exists a numerical constant $C_{1}>0$ such that:
@@
Corollary B.4 refines Theorem 5.1 using a standard peeling argument to substitute the norm $\left\|{x}\right\|$ for the constant radius $r$ in the original probability expression in equation (5.2). This adjustment allows varying distances $x$ from the origin, ranging from $\underline{r}$ to $r$, enhancing the applicability of the probability bounds over different scales. For completeness, we give its proof in Section B.1.
Corollary B.4.
@@
We are now ready to prove Theorem 5.2. We pick
@@
in Corollary B.4. We then derive for some absolute constant $C_{2}$:
@@
We without loss of generality assume that $C_{2}\max\{\Delta_{0},\Delta_{0}^{2}\}\leq\gamma$ where $\gamma$ is the numerical constant specified in Theorem B.2, since otherwise Theorem 5.2 trivially follows from equation (B.1).
@@
In the first case, since $0\in\mathcal{Z}$ by Theorem B.1, we immediately obtain:
@@
Since we have assumed $C_{2}\max\{\Delta_{0},\Delta_{0}^{2}\}\leq\gamma$, by Theorem B.2 we obtain that $x^{*}$ must obey
@@
where $C>0$ is an absolute constant. This completes the proof of Theorem 5.2.
B.1 Proof of Corollary B.4
We can define a sequence $\{r^{(i)}\}_{i=1}^{N}$ so that $r^{(0)}=\underline{r}$, $r^{(N)}=r$, $r^{(i+1)}/r^{(i)}\leq e$ and $N\leq\log(er/\underline{r})$. Under the condition of Corollary B.4, we can apply Theorem 5.1 and obtain for every $1\leq i\leq N$
@@
On this event, for every $x$ with $\underline{r}\leq\left\|{x}\right\|\leq r$, we find $1\leq i(x)\leq N$ such that $r^{(i(x)-1)}\leq x\leq r^{(i(x))}$, and then we obtain that
@@
where the last inequality is due to $r^{(i(x))}/x\leq r^{(i(x))}/r^{(i(x)-1)}\leq e$ by construction. Hence, we have proven the existence of a numerical constant $C^{\prime}>0$ such that:
@@
This completes the proof of Corollary B.4.
Appendix C Proof of Corollary 5.3
We apply Theorem 4.2. We first check the assumptions. For the nonsmooth $h(z)=|z|$, it is decomposed as $h=h^{ns}+h^{sm}$ where $h^{sm}(z)=-z$ and $h^{ns}(z)=2(z)_{+}$.
@@
    where $e(X,\xi)=(h^{sm})^{\prime}(c(X;\xi))\nabla c(X;\xi)$. For every matrix $V\in\mathbb{R}^{D\times D}$, $\langle A,V\rangle$ is $\left\|{V}\right\|_{F}$-subgaussian. Given the property that the sum of two subgaussian random variables is still subgaussian [Ver18, Exercise 2.5.7], we deduce that $\langle A+A^{T},V\rangle$ is $2\left\|{V}\right\|_{F}$-subgaussian for every matrix $V$. As a consequence, we derive that, each of
@@
    must be $2\left\|{X_{1}-X_{2}}\right\|_{F}$ subgaussian, and thus $C\left\|{X_{1}-X_{2}}\right\|_{F}$ subexponential for every matrix $V\in\mathbb{R}^{D\times D}$ with $\left\|{V}\right\|_{F}=1$ for some universal constant $C>0$. Here we use the fact that every sub-gaussian distribution is subexponential [Ver18, Section 2.7],
@@
We then compute the bound in Theorem 4.2. Namely, we need to compute $\zeta$ and ${\rm vc}(\mathcal{F})$.
@@
Note that $X\mapsto\langle A,XX^{T}\rangle-b$ is a degree $2$ polynomial in $X$ for every $A,b$. We then bound its VC dimension using Theorem 6.1, which yields that for some universal constant $C>0$:
@@
Corollary 5.3 then follows by applying Theorem 6.1.

### v2

@@ A.1.2 Bounding Individual Term using VC-Dimension (Proof of Lemma 3)
$\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})\leq\max_{b\in H(\xi_{1:m})}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}b_{i}u_{v}(\xi_{i}).$
@@
$V(\tau,\xi_{1:m})\leq\mathbb{E}_{\varepsilon_{1:m}}\left[\exp\left(\tau\cdot\max_{b\in H(\xi_{1:m})}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}b_{i}u_{v}(\xi_{i})\right)\right].$
@@
$\begin{split}V(\tau,\xi_{1:m})&\leq\mathbb{E}_{\varepsilon_{1:m}}\left[\sum_{b\in H(\xi_{1:m})}\sum_{v\in V_{1/2}}\exp\left(\frac{\tau}{m}\cdot\sum_{i=1}^{m}\varepsilon_{i}b_{i}u_{v}(\xi_{i})\right)\right]\\ &=\sum_{b\in H(\xi_{1:m})}\sum_{v\in V_{1/2}}~{}\prod_{i=1}^{m}\mathbb{E}_{\varepsilon_{i}}\left[\exp\bigg{(}\frac{\tau}{m}\cdot\varepsilon_{i}b_{i}u_{v}(\xi_{i})\bigg{)}\right]\end{split}$ (A.15)
where the symbol $\mathbb{E}_{\varepsilon_{i}}$ means the expectation is taken solely over the Rademacher variables $\varepsilon_{i}$.
Given that for any $b_{i}\in\{0,1\}$ and $q_{i}\in\mathbb{R}$, the expectation over $\varepsilon_{i}$ obeys:
$\mathbb{E}_{\varepsilon_{i}}[\exp(\varepsilon_{i}b_{i}q_{i})]=\cosh(b_{i}q_{i})\leq\cosh(q_{i})=\mathbb{E}_{\varepsilon_{i}}[\exp(\varepsilon_{i}q_{i})]$
@@
$\begin{split}V(\tau,\xi_{1:m})&\leq\sum_{v\in V_{1/2}}\sum_{b\in H(\xi_{1:m})}\prod_{i=1}^{m}\mathbb{E}_{\varepsilon_{i}}\left[\exp\left(\frac{\tau}{m}\cdot\varepsilon_{i}u_{v}(\xi_{i})\right)\right]\\ &=|H(\xi_{1:m})|\sum_{v\in V_{1/2}}\prod_{i=1}^{m}\mathbb{E}_{\varepsilon_{i}}\left[\exp\left(\frac{\tau}{m}\cdot\varepsilon_{i}u_{v}(\xi_{i})\right)\right].\end{split}$
@@
$V(\tau,\xi_{1:m})\leq(2m)^{{\rm vc}(\mathcal{W})}\sum_{v\in V_{1/2}}\prod_{i=1}^{m}\mathbb{E}_{\varepsilon_{i}}\left[\exp\left(\frac{\tau}{m}\cdot\varepsilon_{i}u_{v}(\xi_{i})\right)\right].$ (A.16)
@@
$\begin{split}U(\tau)&\leq\mathbb{E}_{\xi_{1:m}}[V(4\tau,\xi_{1:m})]\leq(2m)^{{\rm vc}(\mathcal{W})}\sum_{v\in V_{1/2}}\prod_{i=1}^{m}\mathbb{E}_{\varepsilon_{i},\xi_{i}}\left[\exp\left(\frac{4\tau}{m}\cdot\varepsilon_{i}u_{v}(\xi_{i})\right)\right]\end{split}$ (A.17)
where $\mathbb{E}_{\varepsilon_{i},\xi_{i}}$ denotes the joint expectation taken over $\varepsilon_{i}$ and $\xi_{i}$.
Since $u_{v}(\xi)$ is $\bar{\sigma}$-sub-exponential, then $\varepsilon_{i}u_{v}(\xi_{i})$ is also $\bar{\sigma}$-sub-exponential for every $v$ by definition. Using [44, Proposition 2.7.1(e)], this implies that, for every $\theta_{2}$, $v$, $\tau\in(0,cm/\bar{\sigma})$:
$\mathbb{E}_{\varepsilon_{i},\xi_{i}}\left[\exp\left(\frac{4\tau}{m}\cdot\varepsilon_{i}u_{v}(\xi_{i})\right)\right]\leq\exp\left(C\frac{\bar{\sigma}^{2}\tau^{2}}{m^{2}}\right)$
@@
By applying Markov’s inequality under the outer integral [31] (which follows from the monotone property of outer integral discussed below Definition 5), we establish that there exists a universal constant $c>0$ such that for every $u\geq 0$:
@@ A.1.3 A Basic Bound
This section proves Lemma 4. By definition:
@@ A.1.4 A Result on Measurability
We prove Lemma 6. Let $\mathcal{Z}_{1}^{o}$ and $\mathbb{R}^{o}$ be any countable dense subsets of $\mathcal{Z}_{1}$ and $\mathbb{R}$, respectively. Our main observation is that the supremum over the uncountable sets $\mathcal{Z}_{1},\mathbb{R}$ is equal to the supremum over the countable sets $\mathcal{Z}_{1}^{o},\mathbb{R}^{o}$, holding for every data instances $\xi_{1:m},\varepsilon_{1:m}$:
$\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})=\sup_{z_{1}\in\mathcal{Z}_{1}^{o},t\in\mathbb{R}^{o}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})$ (A.19)
Since the RHS is the pointwise supremum over countable measurable functions of $\xi_{1:m},\varepsilon_{1:m}$, it is measurable. Therefore, the LHS must also be measurable under the product space. To complete the proof of Lemma 6, it suffices to prove equation (A.19).
Fix the values of $\xi_{1:m}$ and $\varepsilon_{1:m}$. It suffices to show that for every $z_{1}\in\mathcal{Z}_{1}$ and $t\in\mathbb{R}$, there exist $z_{1}^{o}\in\mathcal{Z}_{1}^{o}$ and $t^{o}\in\mathbb{R}^{o}$ such that for every $i=1,2,\ldots,m$:
@@
Recall that $\psi_{z_{1},t}(\xi)=\mathbf{1}\{o(z_{1},\xi)\geq t\}$. First, we note an important observation, which uses the fact that $t\mapsto\mathbf{1}\{w\geq t\}$ is left continuous for every real value $w$. For every $z_{1}$, there exists $t^{o}\in\mathbb{R}^{o}$ with $t^{o}<t$ such that for every $i=1,2,\ldots,m$:
@@
This proves that for every $z_{1}\in\mathcal{Z}_{1}$ and $t\in\mathbb{R}$, there is the existence of the pair $z_{1}^{o}\in\mathcal{Z}_{1}^{o}$ and $t^{o}\in\mathbb{R}^{o}$ such that equation (A.20) holds for all $i=1,2,\ldots,m$. This completes the proof of Lemma 6.
Appendix B Proof of Theorem 7
Our approach comprises two main steps. First, we demonstrate that every stationary point of the empirical objective $\Phi_{S}$ must be approximately stationary for the population objective $\Phi$, leveraging the uniform convergence of subdifferentials in Theorem 6. Second, we utilize established results from the literature [14, Section 5] on these (approximate) stationary points of $\Phi$ to pinpoint the location of empirical stationary points.
To set up the stage, we recall the location of the stationary point and approximate stationary point of the population objective $\Phi$ formally derived in [14].
Theorem 11 ([14, Theorem 5.2]).
@@ Appendix B Proof of Theorem 7
The following Theorem 12 can be easily deduced from [14, Theorem 5.2+Theorem 5.3]. Recall that $\mathcal{Z}=\{x:0\in\partial\Phi(x)\}$ denotes the set of stationary points of $\Phi$.
Theorem 12 ([14, Theorem 5.3]).
@@
${\rm dist}(0;\partial\Phi(x))\leq\gamma\left\|{x}\right\|,$
@@
The following result is immediate given [14, Corollary 6.3].
Proposition 13 ([14, Corollary 6.3]).
There exist a numerical constant $C>0$ so that for $m\geq Cd$, with an inner probability at least $1-2\exp(-d)$ [^1] The original expression in the paper [14], $1-2\exp(-\min\{c_{2}m,d^{2}\})$, should be corrected to $1-2\exp(-d)$. , the set of stationary points $\mathcal{Z}_{S}$ of $\Phi_{S}$ satisfies:
@@
We shall build on top of these existing results to establish Theorem 7. First, Proposition 13 implies there exists a numerical constant $C_{1}>0$ such that:
@@
Corollary 2 refines Theorem 6 using a standard peeling argument to substitute the norm $\left\|{x}\right\|$ for the constant radius $r$ in the original probability expression in equation (5.2). This adjustment allows varying distances $x$ from the origin, ranging from $\underline{r}$ to $r$, enhancing the applicability of the probability bounds over different scales. For completeness, we give its proof in Section B.1.
Corollary 2.
@@
We are now ready to prove Theorem 7. We pick
@@
in Corollary 2. We then derive for some absolute constant $C_{2}$:
@@
We without loss of generality assume that $C_{2}\max\{\Delta_{0},\Delta_{0}^{2}\}\leq\gamma$ where $\gamma$ is the numerical constant specified in Theorem 12, since otherwise Theorem 7 trivially follows from equation (B.1).
@@
In the first case, since $0\in\mathcal{Z}$ by Theorem 11, we immediately obtain:
@@
Since we have assumed $C_{2}\max\{\Delta_{0},\Delta_{0}^{2}\}\leq\gamma$, by Theorem 12 we obtain that $x^{*}$ must obey
@@
where $C>0$ is an absolute constant. This completes the proof of Theorem 7.
B.1 Proof of Corollary 2
We can define a sequence $\{r^{(i)}\}_{i=1}^{N}$ so that $r^{(0)}=\underline{r}$, $r^{(N)}=r$, $r^{(i+1)}/r^{(i)}\leq e$ and $N\leq\log(er/\underline{r})$. Under the condition of Corollary 2, we can apply Theorem 6 and obtain for every $1\leq i\leq N$
@@ B.1 Proof of Corollary 2
On this event, for every $x$ with $\underline{r}\leq\left\|{x}\right\|\leq r$, we find $1\leq i(x)\leq N$ such that $r^{(i(x)-1)}\leq\left\|{x}\right\|\leq r^{(i(x))}$, and then we obtain that
@@
where the last inequality is due to $r^{(i(x))}/\left\|{x}\right\|\leq r^{(i(x))}/r^{(i(x)-1)}\leq e$ by construction. Hence, we have proven the existence of a numerical constant $C^{\prime}>0$ such that:
@@
This completes the proof of Corollary 2.
Appendix C Proof of Corollary 1
We apply Theorem 5. We first check the assumptions. For the nonsmooth $h(z)=|z|$, it is decomposed as $h=h^{ns}+h^{sm}$ where $h^{sm}(z)=-z$ and $h^{ns}(z)=2(z)_{+}$.
@@ Appendix C Proof of Corollary 1
    where $e(X,\xi)=(h^{sm})^{\prime}(c(X;\xi))\nabla c(X;\xi)$. For every matrix $V\in\mathbb{R}^{D\times D}$, $\langle A,V\rangle$ is $\left\|{V}\right\|_{F}$-subgaussian. Given the property that the sum of two subgaussian random variables is still subgaussian [44, Exercise 2.5.7], we deduce that $\langle A+A^{T},V\rangle$ is $C\left\|{V}\right\|_{F}$-subgaussian for every matrix $V$, where $C>0$ is a universal constant. As a consequence, we derive that, each of
@@
    must be $C\left\|{X_{1}-X_{2}}\right\|_{F}$ subgaussian, and thus $C^{\prime}\left\|{X_{1}-X_{2}}\right\|_{F}$ subexponential for every matrix $V\in\mathbb{R}^{D\times D}$ with $\left\|{V}\right\|_{F}=1$ for some universal constant $C^{\prime}>0$. Here we use the fact that every sub-gaussian distribution is subexponential [44, Section 2.7].
  • (Assumption C.3). It is easy to verify that
    $\begin{split}\left\|{\nabla c(X_{1},\xi)-\nabla c(X_{2},\xi)}\right\|_{F}&=\left\|{(A+A^{T})(X_{1}-X_{2})}\right\|_{F}\leq 2\left\|{A}\right\|_{F}\left\|{X_{1}-X_{2}}\right\|_{F}\\ \left\|{e(X_{1},\xi)-e(X_{2},\xi)}\right\|_{F}&=\left\|{(A+A^{T})(X_{2}-X_{1})}\right\|_{F}\leq 2\left\|{A}\right\|_{F}\left\|{X_{1}-X_{2}}\right\|_{F}\end{split},$
    where $e(X,\xi)=(\psi^{sm})^{\prime}(c(X;\xi))\nabla c(X;\xi)$. Since $A$ is $\sigma$-subgaussian, $\mathbb{E}[\left\|{A}\right\|_{F}]<\infty$. As a result, Assumption C.3 is satisfied with $L(\xi)=2\left\|{A}\right\|_{F}$.
@@
We then compute the bound in Theorem 5. Namely, we need to compute $\zeta$ and ${\rm vc}(\mathcal{F})$.
@@
Note that $X\mapsto\langle A,XX^{T}\rangle-b$ is a degree $2$ polynomial in $X$ for every $A,b$. We then bound its VC dimension using Theorem 8, which yields that for some universal constant $C>0$:
@@
Corollary 1 then follows by applying Theorem 8.